OpenStudy (anonymous):

Medal will be Awarded for best answer. What is the exact value of the square root of the quantity 128/14.? Simplify if possible

OpenStudy (anonymous):

\[\frac{ \sqrt{128} }{ \sqrt{14} }\]

OpenStudy (anonymous):

11.31370849898476/3.74165738773941 That is what I got that is the best I can do.

OpenStudy (anonymous):

that's not any of the answers, thanks for trying though

OpenStudy (anonymous):

what are the answers?

OpenStudy (anonymous):

\[\frac{ 8\sqrt{2} }{ \sqrt{2} }\] \[\frac{ 8\sqrt{7} }{ \sqrt{7} }\] \[\frac{ 16\sqrt{2} }{ \sqrt{2}}\] \[\frac{ 16\sqrt{7} }{ \sqrt{7}}\]

jigglypuff314 (jigglypuff314):

ummm... I'm getting (8√7)/7

OpenStudy (anonymous):

Oh well that changes things. Yeah I'm getting (8√7)/7 too

OpenStudy (anonymous):

than I will choose b thanks @jigglypuff314 @Help_PLEASE_18

jigglypuff314 (jigglypuff314):

would you like an explanation of how we got that?

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

yes pls, when I tried this I got A.

jigglypuff314 (jigglypuff314):

√128 = √(64*2) = √64 * √2 = 8√2 so you get (8√2)/√14 but most people don't want roots in the denominator so multiply by (√14/ √14) so (8√2 * √14)/(√14*√14) = (8√2 * √14)/(14) = (8√2*√2*√7)/14 = (8 * 2 √7)/14 = 16√7 / 14 = 8√7 / 7 one must have understood that √14 * √14 = 14 and that √2 * √2 = 2 feel free to ask if you still don't understand :)

OpenStudy (anonymous):

wow thanks, I understand everything!!:)

OpenStudy (anonymous):

my mistake was here: (8√2 * √14)/(√14*√14) = (8√2 * √14)/(14)

jigglypuff314 (jigglypuff314):

I see :) well I'm glad I could help :D